xi^64-3i^108
------------ = 1+i^396
2i^160+i^148.y
FIND x and y.

Answer by stanbon(75887)

Note: All multiple-of-4 powers of i = 1
--------------------------------------------
All of your powers are multiples of 4.
------------------------------------------------
xi^64-3i^108
------------ = 1+i^396
2i^160+i^148.y
FIND x and y.
--------------------------------
(x -3)/(2 + y) = 1 +1
----
(x-3) = 2(2+y)
x-3 = 4+2y
2y-x+7 = 0
===================
You would need another equation in x and y to determine
the values of x and y.
=====================
Cheers,
Stan H.